OpenStudy (anonymous):

integral of (sec^2(8t)tan^2(8t))/(sqrt(16-tan^2(8t)))dt

hartnn (hartnn):

did you try tan 8x = u ?

OpenStudy (anonymous):

I got.. (sec(8t)(16sqrt(34cos(16t)+30))arctan((sqrt(2)sin(8t))/(sqrt(17cos(16t)+15))-(17cos(16t)+15)tan(8t)sec(8t)))/(32sqrt(16-tan^2(8t))) but it says its wrong.. once again.

OpenStudy (anonymous):

i would guess putting \(u=\tan(x), du =\sec^2(x)dx\)

hartnn (hartnn):

your this answer is also correct...i still think there is brackets mismatch somewhere, but couldn't find it :P

OpenStudy (anonymous):

turns in to \[\int\frac{u^2du}{\sqrt{16-u^2}}\]

OpenStudy (anonymous):

lol! okay I'll try to add brackets...

OpenStudy (anonymous):

then another sub \(z=\sin(4x)\) gives \[16\int \sin^2(z)dz\]

OpenStudy (anonymous):

which you can probably look up in a book or else rewrite in terms of \(\cos(2z)\)

hartnn (hartnn):

did you mean z=4sin x ?

OpenStudy (anonymous):

added in brackets and i got it! thanks!

hartnn (hartnn):

welcome ^_^ bdw, you did it on your own, so you deserve a medal, here it is :)